OpenStudy (anonymous):

six stand up comics, A, B, C, D, E, and F, are to perform on a single evening at a comedy club. The order of performance is determined by random selection. Find the probability that: 1. Comic D will perform first. 2. Comic B will perform fifth and comic A will perform last. 3. The comedians will perform in the following order: D, B, A, C, F, E 4. Combic F or combic E will perform third.

OpenStudy (raden):

1) 5!/6! = 5!/6*5! = 1/6

OpenStudy (anonymous):

can you tell me what formula we are using?

OpenStudy (raden):

probability = Number of ways event/ The total number of possible outcomes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

really having a hard time

OpenStudy (raden):

for this case, to find of the even Comic D will perform first, we use permutation concept, 5P5

OpenStudy (anonymous):

how about question 2?

OpenStudy (raden):

2) P=4!/6! = 4!/6(5)(4!) = 1/30

OpenStudy (anonymous):

i was totally off on my answer

OpenStudy (raden):

3) 1/6! = 1/720

OpenStudy (raden):

4) P = 2*5!/6! = 2/6 = 1/3

OpenStudy (anonymous):

okay i'm really lost now!

OpenStudy (raden):

sorry, icant explaind much to u.... maybe u can search it in internet, because many basics that have to u know

OpenStudy (anonymous):

i have the textbook but its not helping me alot.

OpenStudy (anonymous):

thank you so much....

OpenStudy (raden):

welcome..